OpenStudy (anonymous):

What are the roots of the equation x2 + 5x − 3 = 0?

OpenStudy (anonymous):

\(x^2+5x-3=0\) Assuming you copied the question correctly, I would use the quadratic formula to solve this. \[0=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

thank you